OpenStudy (anonymous):

What is the product of five over six and 2? one and two over three one and one over three five over twelve ten over twelve

OpenStudy (jonnyvonny):

When you multiply fractions, all you do is simply multiply across:\[\frac{ 5 }{ 6 }*\frac{ 2 }{1 }=\frac{ 10 }{ 6}\] This, of course can reduce by taking out the values 2 from both the numerator and denominator. This will yield what answer?